Thursday, February 26, 2009

OSG Software Cache, GOC Web Pages and VORS Restored - GOC Ticket # 6442

There was an outage to the OSG Cache, GOC Web Pages, and VORS this evening. Service is now restored. The outage time began at 17:10 EDT and ended at 17:49 EDT.

Wednesday, February 25, 2009

Production BDII server updated (No service interruption)

BDII users,

This morning, we updated various pieces on the central CEMon-BDII service to the latest release that has been under testing on the dev/ITB servers. There was no service interruption - we did the upgrade one server at a time while the other continued to serve users.

Major updates include

1) Use of resource_group name instead of resource in BDII DNs per CMS request. See

2) Resource_alias from OIM is being considered when processing raw cemon LDIF files; For example, NERSC should see their resource's files in the status page.

3) OIM data caching moved out of translation script;
- Caching script also caches Interop BDII LDAP URI list; i.e. [] is cached on both BDII servers []; we are still following up on how WLCG's top level BDII uses this list, and if it is even necessary.

4) Status CGI page [] prints XML for consumption by MyOSG, GIP Validator, etc.

5) BDII upgraded to 4.0.2-2 - no noticeable change functionality but this upgrade allowed us to get rid of a hack we had to put into the BDII update script previously to allow two BDIIs to run on different ports.

Please let us know if you have any questions.

Tuesday, February 24, 2009

GOC services: OIM, MyOSG services restored

> The server that hosts OIM and MyOSG has been down for the last 10
> minutes or so; GOC staff are investigating, and will keep you updated.
> Any systems that depend on OIM data may fail depending on whether they
> cache the OIM data or not.

MyOSG and OIM services have been restored.

The approximate downtime was 35 minutes due to a hardware failure - one that is being investigated further using standard practices (contact hardware vendor, etc.). We apologize for the inconvenience.

Aside: This downtime has lead GOC staff to a better understanding of dependencies between OIM and some other services [most obviously the Interop BDII listing] that should be replicated or cached. We will implement those features, that increase survivability, in the very near future.

Monday, February 23, 2009

New VO Package - Updates to GROW, LIGO, FNAL, and LIGO

There is a new VO package which includes the re-entrance of the GROW VO, a change in address of the LIGO VOMS, updated mappings for the Fermilab VO, and the removal of the SDSS VO.

All OSG resources are encouraged to upgrade. Please remember to backup your OSG installation before proceeding with an update.

For resources that have updated to 0.8.0 and 1.0.0 please use:

pacman -update OSG:vo-package

For OSG Client users please use (Note: This not for wn-client and works for both 0.8 and 1.0.0 releases of the OSG Client.):

pacman -update OSG:vo-client-0.6

Please contact the GOC with any questions.

Thursday, February 19, 2009

Grid Site Administrators Workshop

Here is an announcement of the upcoming Grid Site Administrators Workshop for the OSG Education Coordinator.

I would like to inform you of a focused, hands-on workshop/school for
OSG site administrator :

which will take place at Clemson University March 19-20 (1.5 day course)

There is no registration fee and limited travel support is available from OSG.

The goal of the workshop is to introduce you to the OSG software
stack, its setup, maintenance and operation with the goal of bringing
*your site* up on the Open Science Grid and opening your
science/education to the grid.

We require each participant to have access to *at least one server* on which you can perform the installation.
(pre-course set up will be necessary)

Lectures will be minimal and the focus will be on the hands-on part, where
we expect you to have (root) access to one or two servers at your home

Interested candidates should send an email to
and also apply online (see instructions on the workshop page).

Please submit problems, requests, and questions at:

Thank You,
OSG Grid Operations Center (GOC)
Submit a GOC ticket:
Email/Phone:, 317-278-9699
GOC Homepage:
RSS Feed:

Thursday, February 12, 2009

VDT 1.10.1t has been released


The VDT team is pleased to announce that VDT 1.10.1t has been released.

This release contains updates for:

1. Generic Information Provider - This update fixes a bug in the dCache admin-interface provider.
2. The Gratia probes - Highlights of this update include improved detection of GRAM jobs when using GRAM 4 and bug fixes in reporting PBS/LSF jobs.
3. Gratia-Services and Gratia-Reporting

Further details, including update instructions, are online at:

Please submit problems, requests, and questions at:

Thank You,
OSG Grid Operations Center (GOC)
Submit a GOC ticket:
Email/Phone:, 317-278-9699
GOC Homepage:
RSS Feed:

Monday, February 9, 2009

Welcome to new OSG Operations RSS feed

Welcome! This is the new OSG operations notifications' RSS feed!

If you have not already done so, then please update your feed reader to subscribe to this news feed.

Please submit problems, requests, and questions at:

Tuesday, February 3, 2009


The GOC hosted VOMS was unavailable today, Tuesday, February 3rd, from
12:13 PM to 1:36 PM due to an internal firewall change. The GOC-hosted
VOMS serves the MIS and OSGEDU VOs. This outage did not affect any other OSG VO. The outage was observed during routine monitoring of all OSG VOMS services by the GOC, and corrected.

Please submit problems, requests, and questions at:

Thank You,
OSG Grid Operations Center (GOC)
Submit a GOC ticket:
Email/Phone:, 317-278-9699
GOC Homepage:
RSS Feed: